OpenStudy (anonymous):

graph the following. Identify wheather or not the graph is a funtion. Then evalutute the graph at my specified domain. F(x)= x+5 x<-2 X^2+2x+3 x=>-2 F(x) (3)= F(x) (-4)= F(x) (-2)= Please help me...i will give you a medal if you can help

OpenStudy (anonymous):

Yes, it's a piece-wise function, but still a function. For f(3) you would evaluate it at 3, with whichever equation is at 3. In this case, it's x^2+2x+3. So you plug in 3. 3^2+2(3)+3 = 18 f(-4) you would use x+5 so it would be -4+5=1 f(-2) would be x^2+2x+3 = 3 Hope that helps!

OpenStudy (radar):

use the link provided previously. It worked well for the first one.

OpenStudy (anonymous):

@kaiyne what are the points? to graph it

OpenStudy (anonymous):

Well, value given is the x value and you solve for the y value. So it ends up being (3,18), (-4,1) and (-2,3) Does it make sense how I got those?

OpenStudy (anonymous):

alittle i have one more can i ask you that one?

OpenStudy (anonymous):

Sure, what is it?

OpenStudy (anonymous):

F(x) 2x+1 x =>1 x^2+3 x<1 F(-2)= F(6)= F(1)=

OpenStudy (anonymous):

f(-2) is less than one, so its x^2+3, or (-2)^2+3 = 7. (-2,7) f(6) is greater than one, so you use 2x+1 or 2(6)+1 = 13. (6,13) f(1) is equal to 1, so its still on 2x+1 (since that is x=>1). It ends up being 2(1)+1 = 3. (1,3)

OpenStudy (anonymous):

is it a function?

OpenStudy (anonymous):

thanks so much i get it now :)